The foot of a ladder is placed 6 feet from a wall. If the top of the ladder rests 8 feet up on the wall, how long is the ladder?
Found 2 solutions by solver91311, nerdybill:
Answer by solver91311(24713)


That would depend on whether or not the wall and the ground form a perfect right angle. Based on that presumption, you can proceed one of two ways:

You have a right triangle with legs of 6 and 8, so the hypotenuse, which is the length of the ladder is given by:



Or you could note that 6 is 2 times 3 and 8 is two times 4, so you have a right triangle with legs in the proportion 3:4 which means the hypotenuse must be in proportion 3:4:5, hence 2 times 5 is 10 which is the length of the ladder.



John


Answer by nerdybill(7384)

The foot of a ladder is placed 6 feet from a wall. If the top of the ladder rests 8 feet up on the wall, how long is the ladder?
.
Applying Pythagorean theorem we have:
Let x = length of the ladder
then
x^2 = 6^2 + 8^2
x^2 = 36 + 64
x^2 = 100
x = 10 feet